In the January of the year 2000, I was one more than eleven times as old as my son William. In January of 2009, I was 7 more than 3 times as old as him. How old was my son in January of 2000?
Answer by ankor@dixie-net.com(22740)

Let a = his age in yr 2000
Let w = Will's age in yr 2000
:
In the January of the year 2000, I was one more than eleven times as old as my son William.
a = 11w + 1
:
In January of 2009, I was 7 more than 3 times as old as him.
(nine yrs later)
a + 9 = 3(w+9) + 7
a + 9 = 3w + 27 + 7
a = 3w + 34 - 9
a = 3w + 25
:
Replace a with (11w+1) from the 1st statement
11w + 1 = 3w + 25
11w - 3w = 25 - 1
8w = 24
w =- 24/8
w = 3 yrs old, Will in yr 2000
:
:
Check this, find a
a = 11(3)1
a = 34 yrs old in 2000
Check in the 2nd statement
34 + 9 = 3(3 + 9 ) + 7
43 = 36 + 7